The Petty and Campbell company bond has an annual coupon rate of 6.80%, pays interest semi-annually, and has a face value of $1,000. If there are 22 years to maturity and your required rate of return on bonds of this risk level is 6.00%, what is the bonds value to you? (with this problem solve for its present value and show all work with the formula)

Please show all formulas and work!
Answer #1

Number of periods = n = 22*2 = 44 semiannual periods

Yield I/Y= 6%/2 = 3% or 0.03

Semiannual Payment PMT = 6.80%*1000/2 = $34

Face Value FV = $1000

Hence, PV = PMT[1 - (1+r)-n]/r + FV/(1+r)n = 34(1 - 1.03-44)/0.03 + 1000/1.0344 = $1097.02